Timmyspu:

Find the real world meaning of f(0) and f(5).

Hoodmemes:

honestly what do u think it is?

Timmyspu:

I dont know

JAIDENGYMNASTIC:

The expression f(0) represents the y-intercept on the graph of f(x). The y-intercept of a graph is the point where the graph crosses the y-axis.

darkknight:

f(0) When you have a specific function, f(0) would be the y-intercept, that is what i think it means by the real world meaning. For example if i had a cost function (in dollars) x^2 + 2x + 1 = (x+1)^2, then f(0) = 1,

darkknight:

now f(5) is just what value you get when you set 5 in your function. Like for example if we look back at the cost function , we plug in 5 for x and we find the value (in this scenario dollars) (the dollars). So simply put f(5) is the output of the function f, when 5 is inserted as an input